The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                    http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 51 January 22, 2001

Contents

Abstracts of 8 refereed papers  1
Abstracts of 2 non-refereed papers 6

Abstracts of Refereed Papers

ORFEUS echelle spectra: Molecular hydrogen in disk, IVC, and HVC gas in front of the LMC

Hartmut Bluhm^{1}, Klaas S. de Boer^{1}, Ole Marggraf^{1}, and Philipp Richter^{1,2}

1 Sternwarte, Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 2 Washburn Observatory, University of Wisconsin–Madison, 475 N. Charter Street, Madison, WI 53706, USA

In front of the LMC molecular hydrogen is found in absorption near 0 km s^{−1}, being local disk gas, near +60 km s^{−1} in an intermediate velocity cloud, and near +120 km s^{−1}, being a high velocity halo cloud. The nature of the gas is discussed based on four Orfeus far UV spectra of LMC stars and including data from the ground and from the IUE satellite. The local gas is cool and, given a span of sight lines of only 2.5^{∘}, rather fluffy. The fractional abundance of H_{2} varies from log f = log[N(H_{2})/(2 · N(H_{2}) + N(Hi))] = −5.4 to −3.3. Metal depletions (up to −1.7 dex for Fe) are typical for galactic disk gas. In the intermediate and high velocity gas an apparent underabundance of neutral oxygen points to an ionization level of the gas of about 90 %. H_{2} is detected in intermediate and high velocity gas towards HD 269546. In the intermediate velocity gas we find an H_{2} column density of log(N) ≃ 15.6. The H_{2} excitation indicates that the line of sight samples a cloud at a temperature below 150 K. Column densities are too small to detect the higher UV pumped excitation levels. The high velocity H_{2} (log(N) ≃ 15.6) is highly excited and probably exposed to a strong radiation field. Its excitation temperature exceeds 1000 K. Due to the radial velocity difference between the halo gas and the Milky Way disk, the unattenuated disk radiation is available for H_{2} excitation in the halo. We do not find evidence for an intergalactic origin of this gas; a galactic as well as a Magellanic Cloud origin is possible.

Accepted by: Astronomy & Astrophysics

For preprints, contact hbluhm@astro.uni-bonn.de
Also available from the URL http://arXiv.org/abs/astro-ph/0012238

Circumstellar masers in the Magellanic Clouds

Jacco Th. van Loon^{1}, Albert A. Zijlstra^{2}, Valentín Bujarrabal^{3} and Lars-Åke Nyman^{4,5}

1 Institute of Astronomy, Madingley Road, Cambridge CB3 0HA, United Kingdom 2 UMIST, P.O.Box 88, Manchester M60 1QD, United Kingdom 3 Observatorio Astronómico Nacional, Campus Universitario, Apartado 1143, E-28800 Alcalá de Henares, Spain 4 European Southern Observatory, Casilla 19001, Santiago 19, Chile 5 Onsala Space Observatory, S-439 92 Onsala, Sweden

Results are presented of a search for 22 GHz H_{2}O 6_{16} → 5_{23}, 43 GHz SiO_{v=1}(J = 1 → 0), 86 GHz SiO_{v=1}(J = 2 → 1) and 129 GHz SiO_{v=1}(J = 3 → 2) maser emission from bright IRAS point sources in the Small and Large Magellanic Clouds — mostly circumstellar envelopes around obscured red supergiants and Asymptotic Giant Branch stars (OH/IR stars). The aim of this effort was to test whether the kinematics of the mass loss from these stars depends on metallicity.

H_{2}O maser emission was detected in the red supergiants IRAS04553−6825 and IRAS05280−6910, and tentatively in the luminous IR object IRAS05216−6753 and the AGB star IRAS05329−6708, all in the LMC. SiO_{v=1}(J = 2 → 1) maser emission was detected in IRAS04553−6825.

The double-peaked H_{2}O maser line profiles of IRAS04553−6825 and IRAS05280−6910, in combination with the OH (and SiO) maser line profiles, yield the acceleration of the outflows from these stars. The outflow velocity increases between the H_{2}O masing zône near the dust-formation region and the more distant OH masing zône from v ∼ 18 to 26 km for IRAS04553−6825 and from v ∼ 6 to 17 km s^{−1} for IRAS05280−6910.

The total sample of LMC targets is analysed in comparison with circumstellar masers in the Galactic Centre. The photon fluxes of circumstellar masers in the LMC are found to be very similar to those in the Galactic Centre. The expansion velocities in the LMC appear to be ∼ 20% lower than for similarly bright OH masers in the Galactic Centre, but the data are still consistent with no difference in expansion velocity. OH/IR stars in the LMC appear to have slower accelerating envelopes than OH/IR stars in the Galactic Centre.

The masers in the LMC have blue-asymmetric emission profiles. This may be due to the amplification of stellar and/or free-free radiation, rather than the amplification of dust emission, and may be more pronounced in low metallicity envelopes.

The SiO maser strength increases with the photometric amplitude at 2.2 μm but is independent of the photometric amplitude at 10 μm. This suggests a strong connection between shocks in the dust-free SiO masing zône and the dust formation process. The LMC masers obey the same trend as the Galactic Centre masers.

Appendices describe H_{2}O maser emission from the moderately mass-losing AGB star R Dor in the Milky Way, optical echelle spectroscopy of IRAS04553−6825, and the properties of circumstellar masers in the Galactic Centre.

Accepted by: Astronomy & Astrophysics For preprints, contact jacco@ast.cam.ac.uk

The Optical Gravitational Lensing Experiment. Cepheids in the Magellanic Clouds. VI. Double-Mode Cepheids in the Large Magellanic Cloud

I. Soszyński^{1}, A. Udalski^{1}, M. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1,2}, P. Woźniak^{3},

and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Universidad de Concepción, Departamento de Fisica, Casilla 160–C, Concepción, Chile 3 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present a sample of 76 double-mode Cepheids detected in the 4.5 square degree area in the central part of the LMC. 19 stars from the sample pulsate in the fundamental mode and the first overtone while 57 objects are the first and second overtone pulsators.

We analyze the period ratio of double-mode Cepheids and Fourier parameters of decomposition of the light curves of these objects. We also present location of different type Cepheids from the LMC in the color-magnitude diagram and show the distribution of their V − I color indices.

Accepted by: Acta Astronomica Vol. 50, 451.

For preprints, contact udalski@astrouw.edu.pl
Also available from the URL http://www.arXiv.org/abs/astro-ph/0101341

Optical studies of two LMC X-ray transients: RX J0544.1–7100 and RX J0520.5-6932

M.J. Coe^{1}, I. Negueruela^{2,3}, D.A.H. Buckley ^{4}, N.J. Haigh^{1} & S.G.T.Laycock^{1}

1 Physics and Astronomy Dept., The University, Southampton, SO17 1BJ, UK 2 SAX SDC, Agenzia Spaziale Italiana, c/o Telespazio, via Corcolle 19, 00131 Roma, Italy 3 Observatoire de Strasbourg, 11 rue de l’Universite, Strasbourg 67000, France 4 South African Astronomical Observatory, PO Box 9, Observatory 7935, South Africa

We report observations which confirm the identities of the optical counterpart to the transient sources RX J0544.1-7100 and RX J0520.5-6932. The counterparts are suggested to be a B-type stars. Optical data from the observations carried out at ESO and SAAO, together with results from the OGLE data base, are presented. In addition, X-ray data from the RXTE all-sky monitor are investigated for long term periodicities. A strong suggestion for a binary period of 24.4d is seen in RX J0520.5-6932 from the OGLE data.

Accepted by: Monthly Notices of the Royal Astronomical Society

Also available from the URL http://xxx.soton.ac.uk/abs/astro-ph/0101192

Discovery of a 9-s X-Ray Pulsar, AX J0049-732, in the Small Magellanic Cloud

Masaru Ueno^{1}, Jun Yokogawa^{1}, Kensuke Imanishi^{1}, and Katsuji Koyama^{1}

1 Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto, 606-8502, Japan

We report on an ASCA discovery of a new X-ray pulsar, AX J0049−732, in the Small Magellanic Cloud. The pulse shape is sinusoidal with a barycentric period of 9.1320 ± 0.0004 s. The X-ray spectrum was fitted by an absorbed power-law model with a photon index of 0.6_{−0.7}^{+1.0} and a column density of 1.3_{−1.3}^{+2.9}×10^{22} cm^{−2}. An unabsorbed flux at 0.7–10.0 keV was estimated to be 8×10^{−13} erg cm^{−2} s^{−1} corresponding to an X-ray luminosity of 4 × 10^{35} erg s^{−1} at an Small Magellanic Cloud distance of 62 kpc.

Accepted by: Publications of the Astronomical Society of Japan

For preprints, contact masaru@cr.scphys.kyoto-u.ac.jp

The peculiar cluster HS 327 in the Large Magellanic Cloud: can OH/IR stars and carbon stars be twins?

Jacco Th. van Loon^{1}, A.A. Zijlstra^{2}, L. Kaper^{3}, G.F. Gilmore^{1}, C. Loup^{4}, and J.A.D.L.

Blommaert^{5}

1 Institute of Astronomy, Madingley Road, Cambridge CB3 0HA, United Kingdom 2 UMIST, P.O.Box 88, Manchester M60 1QD, United Kingdom 3 Astronomical Institute, University of Amsterdam, Kruislaan 403, NL-1098 SJ Amsterdam, The Netherlands 4 Institut d’Astrophysique de Paris, 98bis Boulevard Arago, F-75014 Paris, France 5 ISO Data Centre, Astrophysics Div., Science Dept. of ESA, Villafranca del Castillo, P.O.Box 50727, E-28080 Madrid, Spain

The obscured OH/IR star IRAS05298−6957 in the LMC was recently noticed to be member of the small double cluster HS 327 that also contains a carbon star (van Loon et al., 1998, A&A 329, 169). Hence they are coeval and have (nearly) the same progenitor mass, which can only be understood if Hot Bottom Burning (HBB) has prevented IRAS05298−6957 from being a carbon star. We present extensive visual and near-IR photometric data for > 10^{4} stars in and around HS 327, and spectroscopic data for some of the brightest AGB stars amongst these. Colour-magnitude diagrams are used to estimate the age for the cluster and its members, and luminosities are derived for the stars for which spectra have been obtained. The age for IRAS05298−6957 and the carbon star is estimated to be ∼ 200 Myr. This corresponds to a Main-Sequence progenitor mass ∼ 4.0 M_{⊙} — the first direct measurement of the lower mass threshold for HBB. This agrees with stellar evolution models that, however, fail to reproduce the low luminosity of the carbon star.

Accepted by: Astronomy & Astrophysics For preprints, contact jacco@ast.cam.ac.uk

Young star clusters in the Large Magellanic Cloud: NGC 1805 and NGC 1818

R.A.Johnson^{1}, S.F.Beaulieu^{1}, G.F.Gilmore^{1}, J.Hurley^{1}, B.X.Santiago^{2}, N.R.Tanvir^{3},

R.A.W.Elson^{4}

1 Institute of Astronomy, Madingley Road, Cambridge CB3 0HA, United Kingdom 2 Instituto de Fisica, Universidade Federal Rio Grande do Sul, 91510-970 Porto Alegre, RS Brazil 3 Department of Physical Science, University of Hertfordshire, College Lane, Hatfield, AL10 9AB, UK 4 Deceased

We present colour-magnitude diagrams for two rich (≈10^{4}M_{⊙}) Large Magellanic Cloud star clusters with ages ≈10^{7} years, constructed from optical and near-infrared data obtained with the Hubble Space Telescope. These data are part of an HST project to study LMC clusters with a range of ages. In this paper we investigate the massive star content of the young clusters, and determine the cluster ages and metallicities, paying particular attention to Be star and blue straggler populations and evidence of age spreads. We compare our data to detailed stellar population simulations to investigate the turnoff structure of ≈25 Myr stellar systems, highlighting the complexity of the blue straggler phenomenon.

Accepted by: Monthly Notices of the Royal Astronomical Society For preprints, contact raj@ast.cam.ac.uk

Also available from the URL http://www.ast.cam.ac.uk/∼raj/biblio.html

Astrometric imaging of crowded stellar fields with only two SIM pointings

N. Dalal^{1} and K. Griest^{1}

1 Physics Dept., UCSD, La Jolla CA 92093, USA

The Space Interferometry Mission (SIM) will observe sources in crowded fields. Recent work has shown that source crowding can induce significant positional errors in SIM’s astrometric measurements, even for targets many magnitudes brighter than all other crowding sources. Here we investigate whether the spectral decomposition of the fringe pattern may be used to disentangle the overlapping fringes from multiple blended sources, effectively by performing synthesis imaging with two baselines. We find that spectrally dispersed fringes enable SIM to identify and localize a limited number of field sources quite robustly, thereby removing their effect from SIM astrometry and reducing astrometry errors to near photon noise levels. We simulate SIM measurements of the LMC, and show that (a) SIM astrometry will not be corrupted by blending and (b) extremely precise imaging of mildly crowded fields may be performed using only two orthogonal baseline orientations, allowing microarcsecond positional measurements. We lastly illustrate the method’s potential with the example of astrometric microlensing, showing that SIM’s mass and distance measurements of lenses will be untainted by crowding.

Submitted to: The Astrophysical Journal

For preprints, contact endall@astrophys.ucsd.edu
Also available from the URL http://arXiv.org/abs/astro-ph/0101217

Abstracts of Non-Refereed Papers

Stellar Associations and their Field East of LMC 4 in the Large Magellanic Cloud

D. Gouliermis^{1}, K. S. de Boer^{1}, S. C. Keller^{2}, M. Kontizas^{3} and E. Kontizas^{4}

1 Sternwarte der Universität Bonn, Auf dem Hügel 71, D-53121 Bonn, Germany 2 Institute of Geophysics and Planetary Physics, Lawrence Livermore Nat. Lab., L-413, P.O. Box 808, Livermore, CA 94550, U.S.A 3 Department of Astrophysics Astronomy & Mechanics, Faculty of Physics, University of Athens, GR-157 83 Athens, Greece 4 Institute for Astronomy and Astrophysics, National Observatory of Athens, P.O. Box 20048, GR-118 10 Athens, Greece

We report about the stellar content and the luminosity and mass functions of three stellar associations and their field located on the north-east edge of the super-bubble LMC 4 in the Large Magellanic Cloud.

To appear in: “Modes of Star Formation and the Origin of Field Populations”, Heidelberg, Germany, October 2000; to be published in the ASP Conference Series, edited by E. K. Grebel and W. Brandner

For preprints, contact dgoulier@astro.uni-bonn.de
Also available from the URL http://www.astro.uni-bonn.de/~dgoulier/pub.shtml#sect2

Runaway collisions in star clusters

Simon F. Portegies Zwart^{1} and Junichiro Makino^{2} and Stephen L. W. McMillan^{3} and

Piet Hut^{4}

1 Massachusetts Institute of Technology, Cambridge, MA 02139, USA, Hubble Fellow 2 Department of Astronomy, University of Tokyo, 7-3-1 Hongo, Bunkyo-ku,Tokyo 113-0033, Japan 3Department of Physics, Drexel University, Philadelphia, PA 19104, USA 4Institute for Advanced Study, Princeton, NJ 08540, USA

We study the occurrence of physical collisions between stars in young and compact star cluster. The calculations are performed on the GRAPE-4 with the starlab software environment which include the dynamical evolution and the nuclear evolution of all stars and binaries. The selection of the initial conditions is based on existing and well observed star clusters, such as R136 in the 30 Doradus region in the Large Magellanic Cloud and the Arches and Quintuplet star clusters in the vicinity of the Galactic center. Collisions between stars occurred rather frequently in our models. At any time a single star dominates the collision history of the system. The collision rate of this runaway merger scales with the initial relaxation time of the cluster and is independent on other cluster parameters, such as the initial mass function or the initial density profile of the cluster. Subsequent encounters result in a steady grow in mass of the coagulating star, until it escapes or explodes in a supernova. The collision rate in these models is about 2.2 × 10^{−4} collisions per star per Myr for a cluster with an initial relaxation time of 1 Myr.

To appear in: ‘Stellar Collisions, Mergers, and their Consequences’, ASP Conference Series (ed. M. Shara) For preprints, contact spz@space.mit.edu

Also available from the URL http://www.manybody.org